Sulphur trioxide is formed by the reaction between sulphur dioxide and air (more correctly oxygen) in a catalytic reactor, according to the unbalanced equation The reaction conditions give a 50% conversion when stoichiometric quantities of the reactants are used. (a) Balance the equation. (b) Calculate the composition (on a mole fraction basis) of the final mixture leaving the reactor.
Answer #1

The balanced chemical equation is

2SO2 + O2 -----> 2SO3

According to the question, the reaction gives 50% conversion when stoichiometric quantities of the reactants are used. In other words only half of the reactants undergo reaction to give the product.

This means if 2 moles of SO2 and 1 mole of O2 are added, half of them will react to give 1 mole of SO3. while 1 mole SO2 and 0.5 mole O2 will remain unreacted and will leave the reactor.

Thus composition of the final mixture leaving the reactor would be

n(SO2 )= 1 , n(SO3) = 1, n(O2) = 0.5

Total number of moles = 2.5

Mole fraction of SO2 = 1/2.5 = 0.4

Mole fraction of SO3 = 1/2.5 = 0.4

Mole fraction of O2 = 0.5/2.5 = 0.2